The distance an object falls is directly proportional to the square of time it has been falling. After 6 seconds it has fallen 1296 feet. How long will it take to fall 2304 feet?
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

The distance an object falls is directly proportional to the square of time it has been falling. After 6 seconds it has fallen 1296 feet. How long will it take to fall 2304 feet?
...
dalphat^2
d=kt^2
1296 = k*6
1290/6 = k
265=k
..
2304 feet how many seconds?
..
2304 = 265*t
t = 2304/265
t =8.69 seconds.
m.ananth@hotmail.ca

Answer by Alan3354(69443)

The distance an object falls is directly proportional to the square of time it has been falling. After 6 seconds it has fallen 1296 feet. How long will it take to fall 2304 feet?
----------------------
1296 = kt^2 = 36k
k = 36
------------
2304 = kt^2 = 36t^2
t^2 = 2304/36 = 64
t = 8 seconds